OpenStudy (anonymous):

round off each of these measurements to 3 significant figures: 6998 and 1779

OpenStudy (shane_b):

First, convert them to scientific notation as follows: \[6998=6.998 x 10^3\]\[1779=1.779 x 10^3\]Now that you have them in scientific notation you can reduce them to 3 significant figures by rounding as necessary and dropping the last digit: \[6.998 x 10^3=7.00 x10^3\]\[1.779 x 10^3=1.78x10^3\]